OpenStudy (anonymous):

Write in standard form Y+2=7/8(x-3) 7/8 is a fraction btwww. Thank you

OpenStudy (anonymous):

First we will distribute, so y+2=7/8x-21/8

OpenStudy (anonymous):

then just move 7/8x to the left side and 2 to the right side to get -7/8x+y=-37/8